3. In an open economy  A = C + I + G + X – Z, with the assumption that induced imports



exist:



If consumption equals R10 mil., investments R30 mil., government expenditure R45 mil.,



exports R20 mil., imports R5 mil., and the marginal propensity to consume equals 0.7, the



marginal propensity to import 0.25 and the tax rate 0.15:



3.1 What is the size of the autonomous aggregate expenditure (A) in this economy?



3.2 What is the size of the multiplier ()?



3.3 What is the size of equilibrium income (Y0) in this economy?

1
Expert's answer
2021-11-02T10:21:33-0400
"A=10+30+45+20-5=100"

"\mu=\frac{1}{1-0.7}=3.3"

"Y=100\times \frac{1}{1-0.75\times(1-0.15)}=275.86"